OpenStudy (anonymous):

Which statement is true about the product radical 2(5 + radical 8)? It is rational and equal to 7. It is rational and equal to 9. It is irrational and equal to 4 + radical 10. It is irrational and equal to 4 + 5 radical 2.

jimthompson5910 (jim_thompson5910):

What did you get here?

jimthompson5910 (jim_thompson5910):

Were you able to get started at all?

OpenStudy (anonymous):

alright, i am confused on how to do this one. So far I have 5 radical 2 + radical 16?

jimthompson5910 (jim_thompson5910):

so far, so good

jimthompson5910 (jim_thompson5910):

what is the square root of 16?

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

so \[\Large 5\sqrt{2} + \sqrt{16}\] \[\Large 5\sqrt{2} + 4\] \[\Large 4+5\sqrt{2}\] is your answer

jimthompson5910 (jim_thompson5910):

this is irrational since you have a root 2 in there

OpenStudy (anonymous):

Thank you @jim_thompson5910